How is per capita GDP affected by GDP growth and population growth?
Answer #1

Answer : The formula of calculating GDP per capita is following :

GDP per capita = GDP / Population

Now if GDP growth level increase remaining population growth as a constant then the GDP per capita increase.

But if population growth level increase remaining GDP growth as a constant then the GDP per capita decrease.

For example : Let the initial GDP level is 2 and population is 2. So,

GDP per capita = 2 / 2 = 1.

Now let GDP increases and becomes 4 without changing in population. So,

GDP per capita = 4 / 2 = 2.

Here GDP per capita increases.

Again let GDP is constant where population increases and becomes 4. So,

GDP per capita = 2 / 4 = 0.5

Here GDP per capita decreases.

Thus, GDP growth and population growth affect the GDP per capita.